OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to and passes through (2, -6).

OpenStudy (aroub):

that is parallel to .....

OpenStudy (anonymous):

parallel to what?

OpenStudy (lalaly):

sth missin

OpenStudy (anonymous):

ok got that one parallel to and

OpenStudy (anonymous):

now all you have to do is find slope of and good luck : )

OpenStudy (anonymous):

y=-3x+2

OpenStudy (anonymous):

plug it in... (6) = -3(2) + 2

OpenStudy (anonymous):

6=-6+b

OpenStudy (anonymous):

b = 0

OpenStudy (lalaly):

parallel means same slope so y--6=-3(x-2) y+6=-3x+6 y=-3x

OpenStudy (anonymous):

y=-3x

OpenStudy (aroub):

parallel=same slope y=mx+b , m= -3 , (2,-6) -6=-3(2)+b -6=-6+b 0=b y=-3x

OpenStudy (anonymous):

y=6

OpenStudy (anonymous):

tht if parallel 2 x-axcis othrwise we need another info

OpenStudy (anonymous):

lol medal for that :D thanks